A Mountain bar has a mass of 6.9×10−2 kg and a calorie rating of 320 Cal .

What speed would this candy bar have if its kinetic energy were equal to its metabolic energy? [Note: The nutritional calorie, 1 Cal, is equivalent to 1000 calories (1000 cal) as defined in physics. In addition, the conversion factor between calories and joules is as follows: 1 Cal = 1000 cal = 1 kcal = 4186 J.]

Express your answer using two significant figures.
Answer #1

Mass of the bar, m = 6.9 * 10^-2 kg = 0.069 kg

Metabolic energy, E = 320 nutritional calories

= 320 * 1000 cal

= 320 * 4186 J

Kinetic energy = metabolic energy

(1/2) m v^2 = 320 * 4186

v^2 = 320 * 4186 * 2 / 0.069

Hence,

v = 6231.10 m/s = 6.23 km/s